Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn, Page 1 of 12  Not logged in ELOG logo
ID Date Author Group Subject Text Attachments
  224   08 Dec 2024, 02:49 Jörg HallmannMIDShift summarysuccessful measurements of elastomer, gelatine
& bio samples

issue:
  
  223   07 Dec 2024, 02:04 Jörg HallmannMIDShift summarydata collection of bubbles in elastomers
(good progress)

issue: beam loss in the tunnel
  
  222   06 Dec 2024, 07:49 Johannes MoellerMIDStability IssueAnd since the 4th there are some instabilities
in the vacuum levels at M1 and M2 as well.
 Screenshot_from_2024-12-06_07-48-48.png 
  221   06 Dec 2024, 06:37 Johannes MoellerMIDStability IssueThe jumps keep repeating very frequently
since yesterday afternoon. The overall drift
of the beam seems to accalerate as well.
 Screenshot_from_2024-12-06_06-39-34.png 
  220   06 Dec 2024, 02:41 Jörg HallmannMIDShift summary
optimization
of the timing of the involved components
(laser, cameras, FEL)
 2024-12-06Bubble0003.png 
  219   05 Dec 2024, 19:47 Alexey ZozulyaMIDStability IssueToday we have observed multiple large disturbances
to beam stability. These can be characterised
by large jumps in the horizontal beam position,
 2024-12-05_19-43-31.png 
  218   05 Dec 2024, 15:58 Jan GruenertPRCIssueadditional comment to MESSAGE ID: 217

PRC was informed by MID at 8h50
about a new appeared limiation to about 60
  
  217   05 Dec 2024, 09:18 Johannes MoellerMIDIssueMID operation limited to 55 bunches by
SA2_XTD1_OPTICS_PROTECTION/MDL/ATT1 because
of inserted Si 0.5mm. Seems a bit too tight
  
  216   05 Dec 2024, 07:38 Jan GruenertPRCIssueAccelerator module issue

7h20:
BKR/RC inform that there is was issue
with the A12 Klystron filament which had
 JG_2024-12-05_um_07.30.17.png 
  215   05 Dec 2024, 00:06 Jörg HallmannMIDShift summary
machine
runs at 17 keV with about 1.2 mJ

  
  214   04 Dec 2024, 15:34 Jan GruenertPRCStatusBeam delivery Status well after tuning
finished (which went faster than anticipated).
 JG_2024-12-04_um_15.33.09.png 
  213   04 Dec 2024, 07:12 Tomas PopelarSPBShift summaryX-ray delivery

19 keV, ~ 400  - 450 uJ, 1
- 128 pulses, stable delivery
  
  212   01 Dec 2024, 23:07 Tokushi SatoSPBShift summaryX-ray delivery

19 keV, ~450 uJ, 1 - 128 pulses,
stable delivery
  
  211   30 Nov 2024, 23:01 Tokushi SatoSPBShift summaryX-ray delivery

19 keV, ~550 uJ, 1 - 128 pulses,
stable delivery
  
  210   26 Nov 2024, 09:47 Andreas KochXPDSA1 statusConcerning
damage of the SA1 FEL imager, Operation-2024
eLog #206
 SA1_FEL_imager_-_damaged_spot_at_YAG_-_original_screen_position.pngSA1_FEL_imager_-_damaged_spot_at_YAG_-_new_screen_position.png 
  209   24 Nov 2024, 17:23 Jan GruenertPRCStatusStatus

All beamlines in operation.
The attachment ATT#1 shows the beam
delivery of this afternoon, with the pulse
 JG_2024-11-24_um_17.22.58.png 
  208   24 Nov 2024, 14:09 Jan GruenertPRCIssueBeckhoff communication loss SA1-EPS (again)

14h00
Same issue as before, but now EEE cannot
clear the problem anymore remotely.
 JG_2024-11-24_um_16.54.02.png 
  207   24 Nov 2024, 13:47 Jan GruenertPRCIssueBeckhoff communication loss SA1-EPS

12h19:
BKR informs PRC about a limitation to 30
bunches in total for the north branch.
  
  206   23 Nov 2024, 15:11 Jan GruenertPRCIssueDamage on YAG screen of SA1_XTD2_IMGFEL

A
damage spot was detected during the restart
of SA1 beam operation at 14h30.
 JG_2024-11-23_um_14.33.06.pngJG_2024-11-23_um_14.40.23.pngJG_2024-11-23_um_15.14.33.png 
  205   23 Nov 2024, 10:27 Jan GruenertPRCIssueSA1 Beckhoff issue

SA1 blocked
to max. 2 bunches and SA3 limited to max
29 bunches
 JG_2024-11-23_um_12.06.35.pngJG_2024-11-23_um_12.10.23.pngJG_2024-11-23_um_14.33.06.png 
ELOG V3.1.4-7c3fd00